OpenStudy (anonymous):

Choose an equation for the line that is parallel to the given line and passes through the given point for questions 1 and 2. 1. y = 5x + 10; (2, 14) y = x + 4 y = –x – 4 y = 5x – 68 y = 5x + 4 Choose an equation for the line that is parallel to the given line and passes through the given point for questions 1 and 2. 1. y = 5x + 10; (2, 14) y = x + 4 y = –x – 4 y = 5x – 68 y = 5x + 4 @Mathematics

OpenStudy (anonymous):

y= 5(x -14) +2 y=5x -70 + 2 y = 5x -68